OpenStudy (mendicant_bias):

I'm doing a Stokes Theorem problem now, trying to learn/remember all of this stuff quickly. Workings so far posted below momentarily.

OpenStudy (mendicant_bias):

http://i.imgur.com/DFNRdyG.png

OpenStudy (dan815):

ok

OpenStudy (mendicant_bias):

Now just about to calculate \[\triangledown \times F\cdot n \]

OpenStudy (mendicant_bias):

Is n just....n is just also one? Just the k hat vector?

OpenStudy (mendicant_bias):

So \[Curl \ F \cdot n = 2 \cdot 1 = 2\]

OpenStudy (mendicant_bias):

So the integrand is two?

OpenStudy (dan815):

yes

OpenStudy (dan815):

normal line for something in the x y plane points straight up or down in z axis

OpenStudy (dan815):

and nhat is unit normal so, <0,0,-1> or <0,0,1> is a normal line, depending on of u wanna go clockwise or ccw

OpenStudy (mendicant_bias):

Alright, so now the bounds of integration are established by the given curve. Yeah. Going to try to set up the integral now.

OpenStudy (dan815):

ok

OpenStudy (dan815):

u shud use a transformation

OpenStudy (dan815):

make the ellipse into a circle

OpenStudy (dan815):

and do polar

OpenStudy (dan815):

and use jacobian to do transform i know ellipse to circle is straight foward but.. practice some jacobians

OpenStudy (mendicant_bias):

Alright, lemme take a shot at the transform to get the circle in the right shape, I vaguely remember how to do it, going to try now.

OpenStudy (dan815):

ok

OpenStudy (mendicant_bias):

\[x = 2u, \ \ \ y = v\]

OpenStudy (dan815):

yeah

OpenStudy (mendicant_bias):

Wait, is the 2 supposed to be in the denominator, e.g. 2/u? One second.

OpenStudy (dan815):

basially u wanna get into a circle so 4x^2+y^2=4 a^2+b^2=4 so a^2=4x^2 a=2x b^2=y^2 b=y

OpenStudy (mendicant_bias):

\[u^2 + v^2 = 4; \ \ \ r = 2\]

OpenStudy (mendicant_bias):

I need to also take the Jacobian to figure out whatever the integrand is supposed to be multiplied by, trying to remember how to do that properly, heh

OpenStudy (mendicant_bias):

Be back in like 2 minutes, dealing with something IRL

OpenStudy (dan815):

ok

OpenStudy (dan815):

jacobian comes from cross product to get area

OpenStudy (mendicant_bias):

Could you elaborate on that? I've always seen the jacobian as the determinant of a matrix with partial derivatives involved.

OpenStudy (mendicant_bias):

@Jhannybean , could you help me set up this Jacobian? I've forgotten like, half of the stuff on this after putting all my time into ODE...re-reading on it now.

OpenStudy (mendicant_bias):

And do I want\[J(u,v)\]or\[J(x,y)?\]

OpenStudy (mendicant_bias):

(Trying to figure out how to set up a 2x2 determinant matrix in LaTeX atm.)

OpenStudy (jhannybean):

Arg...Jacobians, gotta look this up in my book.

OpenStudy (mendicant_bias):

I know/remember how to do this, but now I'm just trying to figure out how to put it into LaTeX, lul. I'll write out what I'm trying to do "by hand".

OpenStudy (mendicant_bias):

http://i.imgur.com/IupsZ0n.png

OpenStudy (jhannybean):

ohh, `\begin{bmatrix}#&#\\#&# \end{bmatrix}`

OpenStudy (jhannybean):

Just insert your numbers where the pound signs are.

OpenStudy (mendicant_bias):

Alright, cool. I just wrote out a whole page of stuff on my tablet and the program crashed and I lost all of it, the disadvantage of "writing things out" ;_;

OpenStudy (mendicant_bias):

\[\begin{bmatrix}2&0\\0&1 \end{bmatrix}\]

OpenStudy (jhannybean):

WOO!

OpenStudy (jhannybean):

...I need to seriously start on my own calc homework -___- Gotta get HW done and then start reviewing Stokes Theorem since we're beginning to learn it this week!!

OpenStudy (mendicant_bias):

\[\frac {\partial x}{\partial u} = 2 \ du = 2, \ \ \ \frac{\partial y }{\partial v} = du = 1\] Alright, so the integrand, dealing with the Jacobian of the transformation, and the \[\triangledown \times F \cdot n = 2,\]

OpenStudy (mendicant_bias):

"And then start reviewing", lol, you're totally on the ball/ahead, congratulations.

OpenStudy (mendicant_bias):

\[\int\limits_{0}^{2 \pi}\int\limits_{0}^{2}4 r \ dr \ d \theta\]

OpenStudy (mendicant_bias):

(Hesitating on that, looks too simple)

OpenStudy (mendicant_bias):

Which equals 16pi if I did things right.

ganeshie8 (ganeshie8):

it is much simpler than that, you could have avoided Jacobian stuff altogether because the change of variables is linear and uniform

OpenStudy (mendicant_bias):

(wot m8, please elabor8)

OpenStudy (mendicant_bias):

(...ganeshie8)

ganeshie8 (ganeshie8):

\[x = 2u, \ \ \ y = v\] Notice that you're scaling uniformly in x direction by a factor of 2 and in y direction bby a factor of 1. This scaling is same through out the graph. So the area also scales same way : \[dx = 2du, \ \ \ y = dv \implies dxdy = 2dudv\]

OpenStudy (mendicant_bias):

Oh, okay, I thought you meant some wholly alternate method, not just shortcutting the Jacobian, but yeah, that makes sense

ganeshie8 (ganeshie8):

No, I only want to point out that you can avoid that determinant in certain cases. For example, you can avoid it for change of variables of form : x= mu, y = nu dx = mdu dy = ndv the scale factor is simply mn. you don't need to find partials and carry out the process to see this..

ganeshie8 (ganeshie8):

but im sure your prof will cut marks if u dont show him jacobian work

ganeshie8 (ganeshie8):

so you can't use this in your hw problems where your prof expects to see jacobian

OpenStudy (mendicant_bias):

Yuh, well, he never has us turn in homework anyways, which I think is incredibly lame, he just has suggested problems and online, computer-graded stuff.

OpenStudy (mendicant_bias):

But yeah, 16 pi, right?

OpenStudy (mendicant_bias):

Supposed to be 4pi, nevermind, I made a mistake somewhere.

ganeshie8 (ganeshie8):

yeah in change of variables i think

ganeshie8 (ganeshie8):

try \[x = u/2, ~~y = v\]

OpenStudy (mendicant_bias):

Ayyy, I suggested that, too, but I got shot down, lol. I'll try that.

ganeshie8 (ganeshie8):

use the earlier trick : dxdy = 1/2 dudv

OpenStudy (mendicant_bias):

So the Jacobian in this case would just be\[J(u,v)=\frac{1}{2},\]the integrand would be identical, so you would be left with-yep

OpenStudy (mendicant_bias):

The integrand would be one instead of four, which would scale by four identically like the other integral did by four.

OpenStudy (mendicant_bias):

Alright, yeah. Makes sense. Thank you!

ganeshie8 (ganeshie8):

Looks good!